# OneActor: Consistent Subject Generation via Cluster-Conditioned Guidance

### v3

@@ Abstract
Text-to-image diffusion models benefit artists with high-quality image generation. Yet their stochastic nature hinders artists from creating consistent images of the same subject. Existing methods try to tackle this challenge and generate consistent content in various ways. However, they either depend on external restricted data or require expensive tuning of the diffusion model. For this issue, we propose a novel one-shot tuning paradigm, termed as OneActor. It efficiently performs consistent subject generation solely driven by prompts via a learned semantic guidance to bypass the laborious backbone tuning. We lead the way to formalize the objective of consistent subject generation from a clustering perspective, and thus design a cluster-conditioned model. To mitigate the overfitting challenge shared by one-shot tuning pipelines, we augment the tuning with auxiliary samples and devise two inference strategies: semantic interpolation and cluster guidance. These techniques are later verified to significantly enhance the generation quality. Comprehensive experiments show that our method outperforms a variety of baselines with satisfactory subject consistency, superior prompt conformity as well as high image quality. Our method is capable of multi-subject generation and compatible with popular diffusion extensions. Besides, we achieve a 4×4$\times$4 × faster tuning speed than tuning-based baselines and, if desired, avoid increasing inference time. Furthermore, to our best knowledge, we are the first to prove that the semantic space of the diffusion model has the same interpolation property as the latent space does. This property can serve as another promising tool for fine generation control.
@@ 1 Introduction
Diffusion probabilistic models ddpm ; ldm ; dpm1 have shown great success in image generation. As a prominent subset of them, text-to-image (T2I) diffusion models CFG ; sdxl significantly improve the artistic productivity. Designers can simply describe a subject (e.g. a character, an object, an art style) that they desire and then obtain high-quality images of the subject. However, relying on the random sampling, diffusion models fail to maintain a consistent appearance of the subject across the generated images. Taking Fig. 1(a) as an example, ordinary diffusion models denoise a random noise from the noise space to a clean latent code guided by the given prompt, where the clean code corresponds to a salient image in the image space. As shown, if given 4 different prompts of the same subject “hobbit” and random noises, ordinary models will generate 4 hobbits with different identities. In other words, the generation of the ordinary models lacks subject consistency.
@@
Our main contributions are: (1) We revolutionize the consistent subject generation by pioneering the cluster-guided task formalization, which supersedes the laborious backbone tuning with a learned semantic guidance to perform efficient generation. (2) We introduce a novel cluster-conditioned generation paradigm, termed as OneActor, to pursue more generalized, nuanced and controlled consistent subject generation. It highlights an auxiliary augmentation during tuning and features the semantic interpolation and cluster guidance scale strategies. (3) Extensive experiments demonstrate our method’s superior subject consistency, excellent prompt conformity and the 4×4$\times$4 × faster tuning speed without inference time increase. (4) We first establish the interpolation property of the semantic space, offering a promising tool for precise generation control.
2 Related Work
Consistent Text-to-Image Generation. A variety of works aim to overcome the randomness challenge of the diffusion models and generate consistent images of the same subject. To start with, personalization, given several user-provided images of one specific subject, aims to generate consistent images of that subject. Some works ti ; per-use-token2 optimize a semantic token to represent the subject, while others dreambooth ; per-use-entiretuning2 tune the entire diffusion model to learn the distribution of the given subject. Later, per-use-parttuning1 ; per-use-parttuning2 ; per-use-parttuning3 ; per-use-parttuning4 find that tuning partial parameters of the diffusion model is effective enough. Apart from the tuning-based methods above, encoder-based methods per-encoder2 ; per-encoder3 ; per-encoder4 ; IP-Adapter ; elite carefully encode the given subject into a representation and manage to avoid user-tuning. However, depending on external given images, they fail to generate imaginary or novel subjects. Recently, a new consistent subject generation thechosenone task is proposed, which aims to generate images of one subject given only its descriptive prompts. Though, its tuning of the whole diffusion model is expensive and may degrade the generation quality. Later, a new approach consistory design handicraft modules to avoid the tuning process and extend consistent generation from single subject to multiple subjects. Yet the extra modules significantly increase the inference time of every image. Besides, based on the localization of subject, it is incapable of abstract subject generation like artistic style. For this issue, we design a new paradigm from a clustering perspective. Learning only in the semantic space, it requires shorter tuning without necessarily increasing inference time.
Semantic Control of Text-to-Image Generation Models. The semantic control starts from classifier-free guidance CFG , which first proposes a text-conditioned denoising network to perform text-to-image generation. It entangles the semantic and latent space so the prompt can guide the denoising trajectory towards the expected destination. Since then, many works manage to manipulate the semantic space to accomplish various tasks. For personalization, they either learn an extra semantic token to guide the latent to the subject cluster ti or push the token embedding to its core distribution to alleviate the overfitting semantic2 . For image editing, they calculate a residual semantic embedding to indicate the editing direction semantic1 ; semantic3 . These works repeatedly confirm that only manipulating the semantic space is an effective way to harness diffusion models for various goals. Hence, in our novel cluster-guided paradigm, we transform the cluster information into a semantic representation. This representation will later guide the denoising trajectories to the corresponding cluster.
@@ 3 Preliminaries
Once trained, we can sample images $\boldsymbol{x}_{0}$ from a Gaussian noise $\boldsymbol{x}_{t}$ by gradually removing the noise step by step with $\epsilon_{\boldsymbol{\theta}}$. To introduce conditional control, classifier-free guidance CFG trains $\epsilon_{\boldsymbol{\theta}}$ in both unconditional and conditional manners: $\epsilon_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{c}_{\emptyset})$ and $\epsilon_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{c}_{\emptyset})$, where $\boldsymbol{c}$ is the given condition and $\boldsymbol{c}_{\emptyset}$ indicates no condition. Images are then sampled by the combination of two types of outputs:
@@ 4.3 Inference Strategies for Versatile Generation
Figure 3: The observation of semantic interpolation property. We vary the latent guidance scale on the left side and the semantic interpolation scale on the right, respectively. The semantic and latent manipulations show the same effect, which proves our argument.
Proof and Implementation of Semantic Interpolation. Classifier-free guidance CFG proves that the conditional interpolation and extrapolation in the latent space reflect the same linear effect in the image space. We argue that the semantic space of the prompt embbedings also shares this property. To prove it, we carry out a toy experiment on SDXL sdxl . Since empty condition $\boldsymbol{c}_{\emptyset}$ is also a semantic embedding obtained by encoding an empty string, we perform interpolation in semantic space: $\boldsymbol{c}^{\prime}=\boldsymbol{c}_{\emptyset}+g\cdot(\boldsymbol{c}-\boldsymbol{c}_{\emptyset})$. Then we replace $\boldsymbol{c}$ with $\boldsymbol{c}^{\prime}$ in the inference process to generate the images in Fig. 3. As shown, for the latent interpolation (left), when we increase the guidance scale $s$ in Eq. (3) within a proper range, the generated image will be more and more compliant with the prompt. While for the semantic interpolation (right), with $g$ increasing, we can observe the same liner reflection in the image space. It proves that the semantic space also possesses the ability of interpolation. We believe that this is because the semantic and latent space are entangled by the denoising network and thus share some properties in common. With this key insight, we can precisely control the effect of $c_{\Delta}$ by replacing Eq. (8) with:
@@
where $v$ is a controllable semantic scale. We investigate its effect in Appendix A.4. In short, a moderate $v$ leads to a optimal balance of consistency and diversity.
Sample with Cluster-Guided Score. Given a subject-centric prompt $\boldsymbol{p}^{\textit{sub}}$ in the inference stage, we first input its semantic embedding $\boldsymbol{c}^{\textit{sub}}$ to the tuned projector $\boldsymbol{\phi}^{*}$ for two representations: $c_{\Delta}^{\textit{tar}}=\boldsymbol{\phi}^{*}(\boldsymbol{h}^{\textit{tar}},\boldsymbol{c}^{\textit{sub}})$ and $c_{\Delta}^{\textit{aver}}=\frac{1}{N-1}\sum_{i=1}^{N-1}\boldsymbol{\phi}^{*}(\boldsymbol{h}^{\textit{aux}}_{i},\boldsymbol{c}^{\textit{sub}})$. Then we approximate the N-1 auxiliary predictions in Eq. (6) with one prediction under the average condition:
@@
In the second variant, for each $\boldsymbol{p}^{\textit{tar}}_{j}$, we treat it as an individual and perform single-subject tuning for multiple projectors $\boldsymbol{\phi}^{*}_{j}$. Then given multi-subject-centric prompt during inference, we offset the embeddings of multiple base words with the semantic outputs $\mathcal{C}=\{c_{\Delta,j}\}_{j=1}^{L}$ of the projectors. To avoid intersecting leakage of semantic information, we constrain the effect of every $c_{\Delta,j}$ to the subject’s corresponding spatial mask in the denoising process. Since all subjects are independent in this variant, we can continually add new subjects without changing existing ones.
Figure 5: The qualitative comparison between consistent subject generation methods. Though all methods generate consistent images given different prompts, our OneActor refines more details such as the character’s clothes.
@@ 5.2 Qualitative Illustration
Single Subject Generation. We illustrate the visual results of personalization baselines and our OneActor in Fig. 4. As shown, TI generates high-quality and diverse images but fails to maintain the consistency with the target image. With global tuning, DB shows, albeit a certain degree of consistency, limited prompt conformity. IP maintains consistency in some cases, but is unable to adhere closely to the prompt. BL suffers from inferior quality during generation. In contrast, benefiting from the intricate cluster guidance and the undamaged denoising network, our method demonstrates a balance of superior subject consistency, content diversity as well as prompt conformity. We specifically compare our OneActor with other consistent subject generation pipelines. Fig. 5 shows that all pipelines manage to generate consistent and diverse images. Yet our method preserves more consistent details (e.g. the gray shirt of the man, the green coat of the boy).
Multiple Subjects Generation. We show double subjects generation comparison between baselines and the two variants of our OneActor in Fig. 6. Most pipelines, represented by DB and TCO, fail to perform multiple subjects generation. CS extends generation of single subject to multiple subjects. However, CS shows mediocre layout diversity, probably due to their unstable SDSA operations consistory . In comparison, both two variants of our method excellently maintain the appearances of two subjects and display superior image diversity. In general, our method is able to generate different types of subjects including humans, animals and objects (or their combinations). Also, our method is versatile for arbitrary styles and expressions ranging from realistic to imaginary. More qualitative results including triple subjects generation and amusing applications are illustrated in Appendix B and C.
@@ 5.3 Quantitative Evaluation
Metrics. To provide an objective assessment, we carry out comprehensive quantitative experiments. We incorporate multi-objective optimization evaluation and encompasses two objectives, identity consistency and prompt similarity, which are popular in personalization tasks ti ; dreambooth . We instruct ChatGPT chatgpt to generate subject target prompts and templates. For a fair and comprehensive comparison, we adopt two metric settings from the two consistent subject generation pipelines thechosenone ; consistory . For the first setting of thechosenone , the identity consistency is CLIP-I, the normalized cosine similarity between the CLIP clip image embeddings of generated images and that of target image, while the prompt similarity is CLIP-T, the similarity between the CLIP text embeddings of the inference prompts and the CLIP image embeddings of the corresponding generated images. For the second setting of consistory , the identity consistency is the DreamSim score dreamsim which focuses on the foreground subject and the prompt similarity is the CLIP-score clipscore .
@@ 6 Conclusion
This paper proposes a novel one-shot tuning paradigm, OneActor, for consistent subject generation. Leveraging the derived cluster-based score function, we design a cluster-conditioned pipeline that performs a lightweight semantic search without compromising the denoising backbone. During both tuning and inference, we devise several strategies for better performance and efficiency. Extensive and comprehensive experiments demonstrate that intricate semantic guidance is sufficient to maintain superior subject consistency, as achieved by our method. In addition to excellent image quality, prompt conformity and extensibility, our method significantly improves efficiency, with an average tuning time of just 5 minutes and avoidable inference time increase. Notably, the interpolation property of the semantic space proven in this paper is a potential tool for fine generation control.
@@ Appendix C More Qualitative Results
In order to comprehensively showcase the performance of our method, we provide more generation samples covering single subject, double subjects and triple subjects in Figs. 16, 17 and 18.

### v4

@@ Abstract
Text-to-image diffusion models benefit artists with high-quality image generation. Yet their stochastic nature hinders artists from creating consistent images of the same subject. Existing methods try to tackle this challenge and generate consistent content in various ways. However, they either depend on external restricted data or require expensive tuning of the diffusion model. For this issue, we propose a novel one-shot tuning paradigm, termed OneActor. It efficiently performs consistent subject generation solely driven by prompts via a learned semantic guidance to bypass the laborious backbone tuning. We lead the way to formalize the objective of consistent subject generation from a clustering perspective, and thus design a cluster-conditioned model. To mitigate the overfitting challenge shared by one-shot tuning pipelines, we augment the tuning with auxiliary samples and devise two inference strategies: semantic interpolation and cluster guidance. These techniques are later verified to significantly improve the generation quality. Comprehensive experiments show that our method outperforms a variety of baselines with satisfactory subject consistency, superior prompt conformity as well as high image quality. Our method is capable of multi-subject generation and compatible with popular diffusion extensions. Besides, we achieve a 4×4$\times$4 × faster tuning speed than tuning-based baselines and, if desired, avoid increasing the inference time. Furthermore, our method can be naturally utilized to pre-train a consistent subject generation network from scratch, which will implement this research task into more practical applications. Project page: https://johnneywang.github.io/OneActor-webpage/.
footnotetext: Corresponding authors.
@@ 1 Introduction
Diffusion probabilistic models ddpm ; ldm ; diffusion have shown great success in image generation. As a prominent subset of them, text-to-image (T2I) diffusion models CFG ; sdxl significantly improve artistic productivity. Designers can simply describe a subject (e.g. a character, an object, an art style) that they desire and then obtain high-quality images of the subject. However, relying on the random sampling, diffusion models fail to maintain a consistent appearance of the subject across the generated images. Taking Fig. 1(a) as an example, ordinary diffusion models denoise a random noise from the noise space to a clean latent code guided by the given prompt, where the clean code corresponds to a salient image in the image space. As shown, if given 4 different prompts of the same subject “hobbit” and random noises, ordinary models will generate 4 hobbits with different identities. In other words, the generation of the ordinary models lacks subject consistency.
@@
Our main contributions are: (1) We revolutionize the consistent subject generation by pioneering the cluster-guided task formalization, which supersedes the laborious backbone tuning with a learned semantic guidance to perform efficient generation. (2) We introduce a novel cluster-conditioned generation paradigm, termed as OneActor, to pursue more generalized, nuanced and controlled consistent subject generation. It highlights an auxiliary augmentation during tuning and features the semantic interpolation and cluster guidance scale strategies. (3) Extensive experiments demonstrate our method’s superior subject consistency, excellent prompt conformity and the 4×4$\times$4 × faster tuning speed without inference time increase. (4) We first establish the semantic-latent guidance equivalence of T2I models, offering a promising tool for precise generation control.
@@ 2 Preliminaries
Once trained, we can sample images $\boldsymbol{x}_{0}$ from a Gaussian noise $\boldsymbol{x}_{t}$ by gradually removing the noise step by step with $\epsilon_{\boldsymbol{\theta}}$. To introduce conditional control, classifier-free guidance CFG trains $\epsilon_{\boldsymbol{\theta}}$ in both unconditional and conditional manners: $\epsilon_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{c}_{\emptyset})$ and $\epsilon_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{c})$, where $\boldsymbol{c}$ is the given condition and $\boldsymbol{c}_{\emptyset}$ indicates no condition. Images are then sampled by the combination of two types of outputs:
@@ 3.3 Inference Strategies for Versatile Generation
Figure 3: The observation of semantic-latent guidance equivalence property. We vary the latent guidance scale on the left side and the semantic interpolation scale on the right, respectively. The semantic and latent manipulations show the same effect, which proves our argument.
Proof and Implementation of Semantic Interpolation. Classifier-free guidance CFG proves that the conditional interpolation and extrapolation in the latent space, controlled by a guidance scale, reflect the same linear effect in the image space. We argue that the semantic space of the prompt embbedings also shares this property. To prove it, we carry out a toy experiment on SDXL sdxl . Since the empty condition $\boldsymbol{c}_{\emptyset}$ is also a semantic embedding obtained by encoding an empty string, we perform interpolation in the semantic space: $\boldsymbol{c}^{\prime}=\boldsymbol{c}_{\emptyset}+g\cdot(\boldsymbol{c}-\boldsymbol{c}_{\emptyset})$. Then we replace $\boldsymbol{c}$ with $\boldsymbol{c}^{\prime}$ in the inference process to generate the images in Fig. 3. As shown, for the latent interpolation (left), when we increase the guidance scale $s$ in Eq. (3) within a proper range, the generated image will be more and more compliant with the prompt. While for the semantic interpolation (right), with $g$ increasing, we can observe the same liner reflection in the image space. This proves that the semantic space also possesses the ability of guidance interpolation. We believe that this is because the semantic and latent space are entangled by the denoising network and thus share some properties in common. With this key insight, we can precisely control the offset guidance effect of $c_{\Delta}$ by replacing Eq. (8) with:
@@
where $v$ is a controllable semantic scale. We further investigate its effect in Appendix A.4. In short, a moderate $v$ leads to an optimal balance of consistency and diversity.
Sample with Cluster-Guided Score. Given a subject-centric prompt $\boldsymbol{p}^{\textit{sub}}$ in the inference stage, we first input its semantic embedding $\boldsymbol{c}^{\textit{sub}}$ to the tuned projector $\boldsymbol{\phi}^{*}$ for two representations:
$c_{\Delta}^{\textit{tar}}=\boldsymbol{\phi}^{*}(\boldsymbol{h}^{\textit{tar}},\boldsymbol{c}^{\textit{sub}}),$ (16)
$c_{\Delta}^{\textit{aver}}=\frac{1}{N-1}\sum_{i=1}^{N-1}\boldsymbol{\phi}^{*}(\boldsymbol{h}^{\textit{aux}}_{i},\boldsymbol{c}^{\textit{sub}}).$ (17)
Then we approximate the N-1 auxiliary predictions in Eq. (6) with one prediction under the average condition:
@@
In the second variant, for each $\boldsymbol{p}^{\textit{tar}}_{j}$, we treat it as an individual and perform single-subject tuning for multiple projectors $\boldsymbol{\phi}^{*}_{j}$. Then given multi-subject-centric prompt during inference, we offset the embeddings of multiple base words with the semantic outputs $\mathcal{C}=\{c_{\Delta,j}\}_{j=1}^{L}$ of the projectors. To avoid intersecting leakage of semantic information, we constrain the effect of every $c_{\Delta,j}$ to the subject’s corresponding spatial mask in the denoising process. Since all subjects are independent in this variant, we can continually add new subjects without affecting existing ones.
@@ 4.2 Qualitative Illustration
Single Subject Generation. We illustrate the visual results of personalization baselines and our OneActor in Fig. 4. As shown, TI generates high-quality and diverse images but fails to maintain the consistency with the target image. With global tuning, DB shows, albeit a certain degree of consistency, limited prompt conformity. IP maintains consistency in some cases, but is unable to adhere closely to the prompt. BL suffers from inferior quality during generation. In contrast, benefiting from the intricate cluster guidance and the undamaged denoising network, our method demonstrates a balance of superior subject consistency, content diversity as well as prompt conformity. We specifically compare our OneActor with other consistent subject generation pipelines. Fig. 5 shows that all pipelines manage to generate consistent and diverse images. Yet our method preserves more consistent details (e.g. the gray shirt of the man, the green coat of the boy). Additionally, we illustrate more examples in Fig. 6, which reaffirms the excellent performance of our method.
Multiple Subjects Generation. We show double subjects generation comparison between baselines and the two variants of our OneActor in Fig. 7. Most pipelines, represented by DB and TCO, fail to perform multiple subjects generation. CS extends generation of single subject to multiple subjects. However, CS shows mediocre layout diversity, probably due to their unstable SDSA operations consistory . In comparison, both variants of our method excellently maintain the appearances of two subjects and display superior image diversity. In general, our method is able to generate different types of subjects including humans, animals and objects (or their combinations). Also, our method is versatile for arbitrary styles and expressions ranging from realistic to imaginary. More qualitative results including triple subjects generation and amusing applications are illustrated in Appendix B and C.
Figure 5: The qualitative comparison between consistent subject generation methods. Though all methods generate consistent images given different prompts, our OneActor refines more details such as the characters’ clothes.
Figure 6: More qualitative results of OneActor. Our method demonstrates excellent subject consistency across characters, animals and objects.
@@ 4.3 Quantitative Evaluation
Metrics. To provide an objective assessment, we carry out comprehensive quantitative experiments. We incorporate multi-objective optimization evaluation and encompass two objectives, identity consistency and prompt similarity, which are popular in personalization tasks ti ; dreambooth . We instruct ChatGPT chatgpt to generate subject target prompts and templates. For a fair and comprehensive comparison, we adopt two metric settings from the two consistent subject generation pipelines thechosenone ; consistory . For the first setting of thechosenone , the identity consistency is CLIP-I, the normalized cosine similarity between the CLIP clip image embeddings of generated images and that of target image, while the prompt similarity is CLIP-T, the similarity between the CLIP text embeddings of the inference prompts and the CLIP image embeddings of the corresponding generated images. For the second setting of consistory , the identity consistency is the DreamSim score dreamsim which focuses on the foreground subject and the prompt similarity is the CLIP-score clipscore .
@@
5 Related Work
Consistent Text-to-Image Generation. A variety of works aim to overcome the randomness challenge of the diffusion models and generate consistent images of the same subject. To start with, personalization, given several user-provided images of one specific subject, aims to generate consistent images of that subject. Some works ti ; per-use-token2 optimize a semantic token to represent the subject, while others dreambooth ; per-use-entiretuning2 tune the entire diffusion model to learn the distribution of the given subject. Later, per-use-parttuning1 ; per-use-parttuning2 ; per-use-parttuning3 ; per-use-parttuning4 find that tuning partial parameters of the diffusion model is effective enough. Apart from the tuning-based methods above, encoder-based methods per-encoder2 ; per-encoder3 ; per-encoder4 ; IP-Adapter ; elite carefully encode the given subject into a representation and manage to avoid user-tuning. While multi-concept composition customization1 ; customization2 goes a step further to display multiple custom subjects on the same generated image via specially designed attention mechanisms. However, depending on external given images, they fail to generate imaginary or novel subjects. Recently, a new consistent subject generation thechosenone task is proposed, which aims to generate images of one subject given only its descriptive prompts. Though, its tuning of the whole diffusion model is expensive and may degrade the generation quality. Later, a new approach consistory designs handicraft modules to avoid the tuning process and extends consistent generation from single subject to multiple subjects. Yet the extra modules significantly increase the inference time of every image. Besides, based on the localization of the subject, it is incapable of abstract subject generation like artistic style. For these issues, we design a new paradigm from a clustering perspective. Learning only in the semantic space, it requires shorter tuning without necessarily increasing inference time. Compared to the training-free pipelines, our pipeline can be naturally utilized to pre-train a consistent subject generation network from scratch.
Semantic Control of Text-to-Image Generation Models. The semantic control starts from classifier-free guidance CFG , which first proposes a text-conditioned denoising network to perform text-to-image generation. It entangles the semantic and latent space so the prompt can guide the denoising trajectory towards the expected destination. Since then, many works manage to manipulate the semantic space to accomplish various tasks. For personalization, they either learn an extra semantic token to guide the latent to the subject cluster ti or push the token embedding to its core distribution to alleviate the overfitting semantic2 . For image editing, they calculate a residual semantic embedding to indicate the editing direction semantic1 ; semantic3 . Besides, previous works facestudio ; photomaker utilize the semantic interpolation of two conditional embeddings for a mixed visual effect. These works repeatedly confirm that solely manipulating the semantic space is an effective way to harness diffusion models for various goals. Hence, in our novel cluster-guided paradigm, we transform the cluster information into a semantic representation. This representation will later guide the denoising trajectories to the corresponding cluster.
@@ 6 Conclusion
This paper proposes a novel one-shot tuning paradigm, OneActor, for consistent subject generation. Leveraging the derived cluster-based score function, we design a cluster-conditioned pipeline that performs a lightweight semantic search without compromising the denoising backbone. During both tuning and inference, we devise several strategies for better performance and efficiency. Extensive and comprehensive experiments demonstrate that intricate semantic guidance is sufficient to maintain superior subject consistency, as achieved by our method. In addition to excellent image quality, prompt conformity and extensibility, our method significantly improves efficiency, with an average tuning time of just 5 minutes and avoidable inference time increase. Notably, the semantic-latent guidance equivalence property proven in this paper is a potential tool for fine generation control. Furthermore, our method is also feasible to pre-train a consistent subject generation network from scratch, which implements this research task into more practical applications.
@@ Appendix C More Qualitative Results
In order to comprehensively showcase the performance of our method, we provide more generation samples covering single subject, double subjects and triple subjects in Figs. 17, 18 and 19.